# Towards Robust Offline Reinforcement Learning under Diverse Data Corruption

### 19 Jan 2024 v2

@@ 1 Introduction
Most previous studies on offline RL have primarily focused on simulation tasks, where data collection is relatively accurate. However, in real-world environments, data collected by humans or sensors can be subject to random noise or even malicious corruption. For example, during RLHF data collection, annotators may inadvertently or deliberately provide incorrect responses or assign higher rewards for harmful responses. This introduces challenges when applying offline RL algorithms, as constraining the policy to the corrupted data distribution can lead to a significant decrease in performance or a deviation from the original objectives. This limitation can restrict the applicability of offline RL to many real-world scenarios. To the best of our knowledge, the problem of robustly learning an RL policy from corrupted offline datasets remains relatively unexplored. It is important to note that this setting differs from many prior works on robust offline RL (Shi & Chi, 2022; Yang et al., 2022a; Blanchet et al., 2023), which mainly focus on testing-time robustness, i.e., learning from clean data and defending against attacks during evaluation. The most closely related works are Zhang et al. (2022); Wu et al. (2022), which focus on the statistical robustness and stability certification of offline RL under data corruption, respectively. Different from these studies, we aim to develop an offline RL algorithm that is robust to diverse data corruption scenarios.
@@ 2 Related Works
Offline RL. To address the distribution shift problem, offline RL algorithms typically fall into policy constraints-based (Wang et al., 2018; Peng et al., 2019; Fujimoto & Gu, 2021; Kostrikov et al., 2021; Xu et al., 2023), and pessimism-based approaches (Kumar et al., 2020; An et al., 2021; Bai et al., 2022; Ghasemipour et al., 2022; Nikulin et al., 2023). Among these algorithms, ensemble-based approaches (An et al., 2021; Ghasemipour et al., 2022) demonstrate superior performance by estimating the lower-confidence bound (LCB) of Q values for OOD actions. Additionally, weighted imitation-based algorithms (Kostrikov et al., 2021; Xu et al., 2023) enjoy better simplicity and stability compared to pessimism-based approaches.
@@ 3 Preliminaries
Offline RL and IQL. We focus on offline RL, which aims to learn a near-optimal policy from a static dataset $\mathcal{D}$. IQL (Kostrikov et al., 2021) employs expectile regression to learn the value function:
@@ 5.2 Huber Loss for Robust Value Function Learning
Assumption 4 only assumes the existence of the $(1+\nu)$-th moment of the noise $\varepsilon_{s,a,s^{\prime}}$, which is strictly weaker than the standard boundedness or sub-Gaussian assumption. This heavy-tailed assumption is widely used in heavy-tailed bandits (Bubeck et al., 2013; Shao et al., 2018) and heavy-tailed RL (Zhuang & Sui, 2021; Huang et al., 2023). Notably, the second moment of $\varepsilon_{s,a,s^{\prime}}$ may not exist (when $\nu<1$), leading to the square loss inapplicable in the heavy-tailed setting. In contrast, by adopting the Huber regression, we obtain the following theoretical guarantee:
@@ 5.3 Penalizing Corrupted Data via In-dataset Uncertainty
We have noted a key technique within IQL’s implementation - the Clipped Double Q-learning (CDQ) (Fujimoto et al., 2018). CDQ learns two Q functions and utilizes the minimum one for both value and policy updating. To verify its effectiveness, we set another variant of IQL without the CDQ trick, namely “IQL(single Q)”, which only learns a single Q function for IQL updates. As shown in Figure 4 (a) and (b), “IQL(single Q)” notably underperforms IQL on most corruption settings except the dynamics corruption. The intrigue lies with the minimum operator’s correlation to the lower confidence bound (LCB) (An et al., 2021). Essentially, it penalizes corrupted data, as corrupted data typically exhibits greater uncertainty, consequently reducing the influence of these data points on the policy. Further results in Figure 4 (c) and (d) explain why CDQ decreases the performance under dynamics corruption. Since the dynamics corruption introduces heavy-tailedness in the Q target as explained in Section 5.2, the minimum operator would persistently reduce the value estimation, leading to a negative value exploding. Instead, “IQL(single Q)” learns an average Q value, as opposed to the minimum operator, resulting in better performance under dynamics corruption.
@@ 6 Experiments
Experimental Setup. Two parameters, corruption rate $c\in[0,1]$ and corruption scale $\epsilon$ are used to control the cumulative corruption level. Random corruption is applied by adding random noise to the attacked elements of a $c$ portion of the datasets. In contrast, adversarial corruption is introduced through Projected Gradient Descent attack (Madry et al., 2017; Zhang et al., 2020) on the pretrained value functions. A unique case is the adversarial reward corruption, which directly multiplies $-\epsilon$ to the original rewards instead of performing gradient optimization. We utilize the “medium-replay-v2” dataset from (Fu et al., 2020), which better represents real scenarios because it is collected during the training of a SAC agent. The corruption rate is set to $c=0.3$ and the corruption scale is $\epsilon=1$ for the main results. More details about all forms of data corruption used in our experiments are provided in Appendix D.1. We compare RIQL with SOTA offline RL algorithms, including ensemble-based algorithms such as EDAC (An et al., 2021) and MSG (Ghasemipour et al., 2022), as well as ensemble-free baselines like IQL (Kostrikov et al., 2021), CQL (Kumar et al., 2020), and SQL (Xu et al., 2023). To evaluate the performance, we test each trained agent in a clean environment and average their performance over four random seeds. Implementation details and additional empirical results are provided in Appendix D and Appendix E, respectively.
@@ 6.1 Evaluation under Random and Adversarial Corruption
We begin by evaluating various algorithms under random data corruption. Table 1 demonstrates the average normalized performance (calculated as 100×scorerandom scoreexpert scorerandom $100\times\frac{\text{score}-\text{random score}}{\text{expert score}-\text{random score}}$) of different algorithms. Overall, RIQL demonstrates superior performance compared to other baselines, achieving an average score improvement of $37.6\%$ over IQL. In 9 out of 12 settings, RIQL achieves the highest results, particularly excelling in the dynamics corruption settings, where it outperforms other methods by a large margin. Among the baselines, it is noteworthy that BC experiences the least average degradation percentage compared to its score on the clean dataset. This can be attributed to its supervised learning scheme, which avoids incorporating future information but also results in a relatively low score. Besides, offline RL methods like EDAC and MSG exhibit performance degradation exceeding $50\%$, highlighting their unreliability under data corruption.
@@ Acknowledgments
The authors would like to thank Chenlu Ye for insightful discussions and the anonymous reviewers for their valuable feedback.
@@ Appendix A Related Works
Offline RL. Ensuring that the policy distribution remains close to the data distribution is crucial for offline RL, as distributional shift can lead to unreliable estimation (Levine et al., 2020). Offline RL algorithms typically fall into two categories to address this: those that enforce policy constraints on the learned policy (Wang et al., 2018; Fujimoto et al., 2019; Peng et al., 2019; Fujimoto & Gu, 2021; Kostrikov et al., 2021; Chen et al., 2021; Emmons et al., 2021; Xu et al., 2023), and those that learn pessimistic value functions to penalize OOD actions (Kumar et al., 2020; Yu et al., 2020; An et al., 2021; Bai et al., 2022; Yang et al., 2022a; Ghasemipour et al., 2022; Sun et al., 2022; Nikulin et al., 2023). Among these algorithms, ensemble-based approaches (An et al., 2021; Ghasemipour et al., 2022) demonstrate superior performance by estimating the lower-confidence bound (LCB) of Q values for OOD actions based on uncertainty estimation. Additionally, imitation-based (Emmons et al., 2021) or weighted imitation-based algorithms (Wang et al., 2018; Nair et al., 2020; Kostrikov et al., 2021) generally enjoy better simplicity and stability compared to other pessimism-based approaches. Studies in offline goal-conditioned RL (Yang et al., 2023; 2022b) further demonstrate that weighted imitation learning offers improved generalization compared to pessimism-based offline RL methods. In addition to traditional offline RL methods, recent studies employ advanced techniques such as transformer (Chen et al., 2021; Chebotar et al., 2023; Yamagata et al., 2023) and diffusion models (Janner et al., 2022; Hansen-Estruch et al., 2023; Wang et al., 2022) for sequence modeling or function class enhancement, thereby increasing the potential of offline RL to tackle more challenging tasks.
@@
Robust RL. One type of robust RL is the distributionally robust RL, which aims to learn a policy that optimizes the worst-case performance across MDPs within an uncertainty set, typically framed as a Robust MDP problem (Nilim & Ghaoui, 2003; Iyengar, 2005; Ho et al., 2018; Moos et al., 2022). Recently, Hu et al. (2022) prove that distributionally robust RL can effectively reduce the sim-to-real gap. Besides, numerous studies in the online setting have explored robustness to perturbations on observations (Zhang et al., 2020; 2021), actions (Pinto et al., 2017; Tessler et al., 2019), rewards (Wang et al., 2020; Husain et al., 2021), and dynamics (Mankowitz et al., 2019). There is also a line of theory works (Lykouris et al., 2021; Wu et al., 2021; Wei et al., 2022; Ye et al., 2023a) studying online corruption-robust RL. In the offline setting, a number of works focus on testing-time robustness or distributional robustness in offline RL (Zhou et al., 2021; Shi & Chi, 2022; Hu et al., 2022; Yang et al., 2022a; Panaganti et al., 2022; Blanchet et al., 2023). Regarding the training-time robustness of offline RL, Li et al. (2023) investigated reward attacks in offline RL, revealing that certain dataset biases can implicitly enhance offline RL’s resilience to reward corruption. Wu et al. (2022) propose a certification framework designed to ascertain the number of tolerable poisoning trajectories in relation to various certification criteria. From a theoretical perspective, Zhang et al. (2022) studied offline RL under contaminated data. One concurrent work (Ye et al., 2023b) leverages uncertainty weighting to tackle reward and dynamics corruption with theoretical guarantees. Different from this purely theoretical work, we propose an algorithm that is both provable and practical under diverse data corruption on all elements.
@@ D.2 Implementation Details of IQL and RIQL
For the policy and value networks of IQL and RIQL, we utilize an MLP with 2 hidden layers, each consisting of 256 units, and ReLU activations. These neural networks are updated using the Adam optimizer with a learning rate of $3\times 10^{-4}$. We set the discount factor as $\gamma=0.99$, the target networks are updated with a smoothing factor of 0.005 for soft updates. The hyperparameter $\beta$ and $\tau$ for IQL and RIQL are set to 3.0 and 0.7 across all experiments. In terms of policy parameterization, we argue that RIQL is robust to different policy parameterizations, such as deterministic policy and diagonal Gaussian policy. For the main results, IQL and RIQL employ a deterministic policy, which means that maximizing the weighted log-likelihood is equivalent to minimizing a weighted $l_{2}$ loss on the policy output: $\mathcal{L}_{\pi}(\phi)=\mathbb{E}_{(s,a)\sim\mathcal{D}}[\exp(\beta A(s,a))\|a-\pi_{\phi}(s)\|_{2}^{2}]$. We also include a discussion about the diagonal Gaussian parameterization in Appendix E.4. In the training phase, we train IQL, RIQL, and other baselines for $3\times 10^{6}$ steps following (An et al., 2021), which corresponds to 3000 epochs with 1000 steps per epoch. The training is performed using a batch size of 256. For evaluation, we rollout each agent in the clean environment for 10 trajectories (maximum length equals 1000) and average the returns. All reported results are averaged over four random seeds. As for the specific hyperparameters of RIQL, we search $K\in\{3,\textbf{5}\}$ and quantile $\alpha\in\{\textbf{0.1},\textbf{0.25},0.5\}$ for the quantile Q estimator, and $\delta\in\{\textbf{0.1},0.5,\textbf{1.0},3.0\}$ for the Huber loss. In most settings, we find that the highlighted hyperparameters often yield the best results. The specific hyperparameters used for the random and adversarial corruption experiment in Section 6.1 are listed in Table 3 and Table 4, respectively. Our anonymized code is available at https://github.com/YangRui2015/RIQL, which is based on the open-source library of (Tarasov et al., 2022).
@@ Appendix E Additional Experiments
Figure 8: Ablation results of IQL under mixed attack.

### 09 Mar 2024 v3

@@ 1 Introduction
Most previous studies on offline RL have primarily focused on simulation tasks, where data collection is relatively accurate. However, in real-world environments, data collected by humans or sensors can be subject to random noise or even malicious corruption. For example, during RLHF data collection, annotators may inadvertently or deliberately provide incorrect responses or assign higher rewards for harmful responses. This introduces challenges when applying offline RL algorithms, as constraining the policy to the corrupted data distribution can lead to a significant decrease in performance or a deviation from the original objectives. This limitation can restrict the applicability of offline RL to many real-world scenarios (Wang et al., 2024; Sun, 2023). To the best of our knowledge, the problem of robustly learning an RL policy from corrupted offline datasets remains relatively unexplored. It is important to note that this setting differs from many prior works on robust offline RL (Shi & Chi, 2022; Yang et al., 2022a; Blanchet et al., 2023), which mainly focus on testing-time robustness, i.e., learning from clean data and defending against attacks during evaluation. The most closely related works are Zhang et al. (2022); Ye et al. (2023b); Wu et al. (2022), which focus on the statistical robustness and stability certification of offline RL under data corruption, respectively. Different from these studies, we aim to develop an offline RL algorithm that is robust to diverse data corruption scenarios.
@@ 2 Related Works
Offline RL. To address the distribution shift problem, offline RL algorithms typically fall into policy constraints-based (Wang et al., 2018; Peng et al., 2019; Fujimoto & Gu, 2021; Kostrikov et al., 2021; Sun et al., 2024; Xu et al., 2023), and pessimism-based approaches (Kumar et al., 2020; An et al., 2021; Bai et al., 2022; Ghasemipour et al., 2022; Nikulin et al., 2023). Among these algorithms, ensemble-based approaches (An et al., 2021; Ghasemipour et al., 2022) demonstrate superior performance by estimating the lower-confidence bound (LCB) of Q values for OOD actions. Additionally, weighted imitation-based algorithms (Kostrikov et al., 2021; Xu et al., 2023) enjoy better simplicity and stability compared to pessimism-based approaches.
@@ 3 Preliminaries
Offline RL and IQL. We focus on offline RL, which aims to learn a near-optimal policy from a static dataset $\mathcal{D}$. IQL (Kostrikov et al., 2021) employs expectile regression to learn the value functions:
@@ 5.2 Huber Loss for Robust Value Function Learning
Assumption 4 only assumes the existence of the $(1+\nu)$-th moment of the noise $\varepsilon_{s,a,s^{\prime}}$, which is strictly weaker than the standard boundedness or sub-Gaussian assumption. This heavy-tailed assumption is widely used in heavy-tailed bandits (Bubeck et al., 2013; Shao et al., 2018) and heavy-tailed RL (Zhuang & Sui, 2021; Huang et al., 2023). Notably, the second moment of $\varepsilon_{s,a,s^{\prime}}$ may not exist when $\nu<1$, leading to the square loss inapplicable in the heavy-tailed setting. In contrast, by adopting the Huber regression, we obtain the following theoretical guarantee:
@@ 5.3 Penalizing Corrupted Data via In-dataset Uncertainty
We have noted a key technique within IQL’s implementation - the Clipped Double Q-learning (CDQ) (Fujimoto et al., 2018). CDQ learns two Q functions and utilizes the minimum one for both value and policy updating. To verify its effectiveness, we set another variant of IQL without the CDQ trick, namely “IQL(single Q)”, which only learns a single Q function for IQL updates. As shown in Figure 4 (a) and (b), “IQL(single Q)” notably underperforms IQL on most corruption settings except the dynamics corruption. The intrigue lies within the minimum operator’s correlation to the lower confidence bound (LCB) (An et al., 2021). Essentially, it penalizes corrupted data, as corrupted data typically exhibits greater uncertainty, consequently reducing the influence of these data points on the policy. Further results in Figure 4 (c) and (d) explain why CDQ decreases the performance under dynamics corruption. Since the dynamics corruption introduces heavy-tailedness in the Q target as explained in Section 5.2, the minimum operator would persistently reduce the value estimation, leading to a negative value exploding. Instead, “IQL(single Q)” learns an average Q value, as opposed to the minimum operator, resulting in better performance under dynamics corruption.
@@ 6 Experiments
Experimental Setup. Two parameters, corruption rate $c\in[0,1]$ and corruption scale $\epsilon$ are used to control the cumulative corruption level. Random corruption is applied by adding random noise to the attacked elements of a $c$ portion of the datasets. In contrast, adversarial corruption is introduced through Projected Gradient Descent attack (Madry et al., 2017; Zhang et al., 2020) with the pretrained value functions. A unique case is the adversarial reward corruption, which directly multiplies $-\epsilon$ to the original rewards instead of performing gradient optimization. We utilize the “medium-replay-v2” dataset from (Fu et al., 2020), which better represents real scenarios because it is collected during the training of a SAC agent. The corruption rate is set to $c=0.3$ and the corruption scale is $\epsilon=1$ for the main results. More details about all forms of data corruption used in our experiments are provided in Appendix D.1. We compare RIQL with SOTA offline RL algorithms, including ensemble-based algorithms such as EDAC (An et al., 2021) and MSG (Ghasemipour et al., 2022), as well as ensemble-free baselines like IQL (Kostrikov et al., 2021), CQL (Kumar et al., 2020), and SQL (Xu et al., 2023). To evaluate the performance, we test each trained agent in a clean environment and average their performance over four random seeds. Implementation details and additional empirical results are provided in Appendix D and Appendix E, respectively.
@@ 6.1 Evaluation under Random and Adversarial Corruption
We begin by evaluating various algorithms under random data corruption. Table 1 demonstrates the average normalized performance (calculated as 100×scorerandom scoreexpert scorerandom $100\times\frac{\text{score}-\text{random score}}{\text{expert score}-\text{random score}}$) of different algorithms. Overall, RIQL demonstrates superior performance compared to other baselines, achieving an average score improvement of $37.6\%$ over IQL. In 9 out of 12 settings, RIQL achieves the highest results, particularly excelling in the dynamics corruption settings, where it outperforms other methods by a large margin. Among the baselines, it is noteworthy that BC experiences the least average degradation percentage compared to its score on the clean dataset. This can be attributed to its supervised learning scheme, which avoids incorporating future information but also results in a relatively low score. Besides, offline RL methods like EDAC and MSG exhibit performance degradation exceeding $60\%$, highlighting their unreliability under data corruption.
@@ Acknowledgments
The work is done when Rui Yang and Tong Zhang were at HKUST. The authors would like to thank Chenlu Ye for insightful discussions and the anonymous reviewers for their valuable feedback.
@@ Appendix A Related Works
Offline RL. Ensuring that the policy distribution remains close to the data distribution is crucial for offline RL, as distributional shift can lead to unreliable estimation (Levine et al., 2020). Offline RL algorithms typically fall into two categories to address this: those that enforce policy constraints on the learned policy (Wang et al., 2018; Fujimoto et al., 2019; Peng et al., 2019; Li et al., 2020; Fujimoto & Gu, 2021; Kostrikov et al., 2021; Chen et al., 2021; Emmons et al., 2021; Sun et al., 2024; Xu et al., 2023), and those that learn pessimistic value functions to penalize OOD actions (Kumar et al., 2020; Yu et al., 2020; An et al., 2021; Bai et al., 2022; Yang et al., 2022a; Ghasemipour et al., 2022; Sun et al., 2022a; Nikulin et al., 2023). Among these algorithms, ensemble-based approaches (An et al., 2021; Ghasemipour et al., 2022) demonstrate superior performance by estimating the lower-confidence bound (LCB) of Q values for OOD actions based on uncertainty estimation (Loquercio et al., 2020; Sun et al., 2022b). Additionally, imitation-based (Emmons et al., 2021) or weighted imitation-based algorithms (Wang et al., 2018; Nair et al., 2020; Kostrikov et al., 2021) generally enjoy better simplicity and stability compared to other pessimism-based approaches. Studies in offline goal-conditioned RL (Yang et al., 2023; 2022b) further demonstrate that weighted imitation learning offers improved generalization compared to pessimism-based offline RL methods. In addition to traditional offline RL methods, recent studies employ advanced techniques such as transformer (Chen et al., 2021; Chebotar et al., 2023; Yamagata et al., 2023) and diffusion models (Janner et al., 2022; Hansen-Estruch et al., 2023; Wang et al., 2022) for sequence modeling or function class enhancement, thereby increasing the potential of offline RL to tackle more challenging tasks.
@@
Robust RL. One type of robust RL is the distributionally robust RL, which aims to learn a policy that optimizes the worst-case performance across MDPs within an uncertainty set, typically framed as a Robust MDP problem (Nilim & Ghaoui, 2003; Iyengar, 2005; Ho et al., 2018; Moos et al., 2022). Hu et al. (2022) prove that distributionally robust RL can effectively reduce the sim-to-real gap. Besides, numerous studies in the online setting have explored robustness to perturbations on observations (Zhang et al., 2020; 2021), actions (Pinto et al., 2017; Tessler et al., 2019), rewards (Wang et al., 2020; Husain et al., 2021), and dynamics (Mankowitz et al., 2019). There is also a line of theory works (Lykouris et al., 2021; Wu et al., 2021; Wei et al., 2022; Ye et al., 2023a) studying online corruption-robust RL. In the offline setting, a number of works focus on testing-time robustness or distributional robustness in offline RL (Zhou et al., 2021; Shi & Chi, 2022; Hu et al., 2022; Yang et al., 2022a; Panaganti et al., 2022; Wen et al., 2023; Blanchet et al., 2023). Regarding the training-time robustness of offline RL, Li et al. (2023) investigated reward attacks in offline RL, revealing that certain dataset biases can implicitly enhance offline RL’s resilience to reward corruption. Wu et al. (2022) propose a certification framework designed to ascertain the number of tolerable poisoning trajectories in relation to various certification criteria. From a theoretical perspective, Zhang et al. (2022) studied offline RL under contaminated data. One concurrent work (Ye et al., 2023b) leverages uncertainty weighting to tackle reward and dynamics corruption with theoretical guarantees. Different from prior work, we propose an algorithm that is both provable and practical under diverse data corruption on all elements.
@@ D.2 Implementation Details of IQL and RIQL
For the policy and value networks of IQL and RIQL, we utilize an MLP with 2 hidden layers, each consisting of 256 units, and ReLU activations. These neural networks are updated using the Adam optimizer with a learning rate of $3\times 10^{-4}$. We set the discount factor as $\gamma=0.99$, the target networks are updated with a smoothing factor of 0.005 for soft updates. The hyperparameter $\beta$ and $\tau$ for IQL and RIQL are set to 3.0 and 0.7 across all experiments. In terms of policy parameterization, we argue that RIQL is robust to different policy parameterizations, such as deterministic policy and diagonal Gaussian policy. For the main results, IQL and RIQL employ a deterministic policy, which means that maximizing the weighted log-likelihood is equivalent to minimizing a weighted $l_{2}$ loss on the policy output: $\mathcal{L}_{\pi}(\phi)=\mathbb{E}_{(s,a)\sim\mathcal{D}}[\exp(\beta A(s,a))\|a-\pi_{\phi}(s)\|_{2}^{2}]$. We also include a discussion about the diagonal Gaussian parameterization in Appendix E.4. In the training phase, we train IQL, RIQL, and other baselines for $3\times 10^{6}$ steps following (An et al., 2021), which corresponds to 3000 epochs with 1000 steps per epoch. The training is performed using a batch size of 256. For evaluation, we rollout each agent in the clean environment for 10 trajectories (maximum length equals 1000) and average the returns. All reported results are averaged over four random seeds. As for the specific hyperparameters of RIQL, we search $K\in\{3,\textbf{5}\}$ and quantile $\alpha\in\{\textbf{0.1},\textbf{0.25},0.5\}$ for the quantile Q estimator, and $\delta\in\{\textbf{0.1},0.5,\textbf{1.0},3.0\}$ for the Huber loss. In most settings, we find that the highlighted hyperparameters often yield the best results. The specific hyperparameters used for the random and adversarial corruption experiment in Section 6.1 are listed in Table 3 and Table 4, respectively. Our code is available at https://github.com/YangRui2015/RIQL, which is based on the open-source library of CORL (Tarasov et al., 2022).
@@ E.1 Ablation of IQL
Figure 8: Ablation results of IQL under mixed attack.